OpenStudy (anonymous):

Solve Each system of equations with substitution 5a-v=-23 -a+v=3

OpenStudy (anonymous):

is this saying that a and v = 3

OpenStudy (thebatman):

-a.

OpenStudy (theslytherinhelper):

All you have to do is pick your favorite variable from one equation (v or a) and solve for it, then substitute whatever answer you get into the other equation.

OpenStudy (anonymous):

solve for what

OpenStudy (anonymous):

idk please just help me solve it

OpenStudy (theslytherinhelper):

I think the 'v' variable would be easiest to work on here

OpenStudy (anonymous):

if u want to solve for a u plug in v-3 into the top equation for a. if u want to solve for v u plug in 3+a into the top equation for v

OpenStudy (theslytherinhelper):

5a-v=-23 -a+v=3 (copied so that I may see it properly) 5a-v=-23 Subtract 5a from both sides now -v = -23 -5a But the variable cannot be negative. Multiply both sides by -1. v = 23 + 5a Now substitute this equation into - a + v = 3, in the 'v' slot.